Suppose you travel North for 35 kilometers, then travel East 65 kilometers. How far are you from your starting point? North and East can be considered the directions of the y and x axis respectively. Round to the tenth's place.
Anser:
Show your work in this space.
to me this seems kinda simple. I have tried 35 + 64 = 100 but I don't think that is the correct anser. The reason for this is if you put it in a graph and you draw a straight line from the starting point to the ending point, then the answer would be 65. I'm just not sure how this question is ment.

Answer by rmromero(383)


Suppose you travel North for 35 kilometers, then travel East 65 kilometers. How far are you from your starting point? North and East can be considered the directions of the y and x axis respectively. Round to the tenth's place.
What is asked in the problem?
How far are you from the starting point?
Given.
You travel north for 35 kilometers
Then travel east 65 kilometers.

Illustration
Travel 65 km East
(Leg)
_____
| /
| /
Tavel 35 km north | / <- this is the distance from the starting point
(Leg) | / (Hypotenuse)
|/
. starting point

Notice that you form a right triangle. we use pythagorean thereom
Pythagorean Theorem states tht in a right triangle, the square of the length
of the hypotenuse is equal to the sum of the squares of the lengths of the legs
See illustration above.

Representation
Let c = the Hypotenuse
a = the leg (north) = 35km
b = the leg (east) = 65
c^2 = a^2 + b^2
c^2 = 35^2 + 65^2
c^2 = 1225 + 4225
c^2 = 5450 extract the sqrt both sides to find c.
sqrt c^2 = sqrt 5450
c = 73.8 km <-------your distance from the starting point